OpenStudy (anonymous):

"t is a real number and p = (x,y) is the point on the unit circle that corresponds to t. find the exact values of the six trigonometric functions of t." (square root 3/2, 1/2) t would be pi/6, right? how would you solve this problem?

OpenStudy (anonymous):

help omg

OpenStudy (anonymous):

ELP PLS

OpenStudy (dumbcow):

you are correct, angle is pi/6 cos(pi/6) = sqrt3/2 sin(pi/6) = 1/2 tan = sin/cos sec = 1/cos csc = 1/sin cot = 1/tan

OpenStudy (anonymous):

i don't understand how to do that, how do i find sec csc and cot?

OpenStudy (anonymous):

sin(pi/6) = 1/2 ; cos(pi/6) = (square root 3/2) ; tan(pi/6) = 1/square root 3 over 2/2 ????